OpenStudy (anonymous):

So 1 to any power is just like multiplying one to the said number power?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Not exactly. It means if you have 1 to the power of something, you would multiply the number 1 times itself as many times as the exponent wants. For example, 1^5 = 1 * 1 * 1 * 1 * 1 = 1.

OpenStudy (anonymous):

@VendettaSoup Thanks! Now I understand that a bit better.

OpenStudy (anonymous):

Anytime!